Sales: $250,000

Net Cost of Purchases: $70,000

Ending Inventory: $ 20,000

Gross Margin: $80,000

Net Income/Loss: $ 24,000

Calculate Beginning Inventory, COGS and Operating Expenses.
Answer #1

COGS = Sales - Gross margin

= 250000 - 80000

= 170000

COGS = Beginning inventory + Net cost of purchases - Ending inventory

170000 = Beginning inventory + 70000 - 20000

Beginning inventory = 170000 - 70000 + 20000

= 120000

Operating expenses = Gross margin - Net income

= 80000 - 24000

= 56000